OpenStudy (anonymous):

Can somebody help me find an expression for the series 2, 2.5, 2.75, 2.875, ... ?

OpenStudy (amistre64):

+0, +1/2, +3/4, + 5/8 maybe? id have to dbl chk that

OpenStudy (amistre64):

nah, that 2.625 in the end .... bummer

OpenStudy (amistre64):

7/8 on the end

OpenStudy (amistre64):

looks like the underside is going by powers of 2 tho

OpenStudy (amistre64):

the top, 0,1,3,7 ; not alot to work off of

OpenStudy (amistre64):

1/2 = .5 1/2 + 1/4 = .75 1/2 + 1/4 + 1/8 = .875

OpenStudy (amistre64):

An = A{n-1} + 1/2^n

OpenStudy (anonymous):

\[\left \{2+\frac{2^n-1}{2^n}:n\in\mathbb{N}\right \}\]

OpenStudy (amistre64):

do we start at n=0 or n=1?

OpenStudy (amistre64):

A1 = 1 + 1/2^(1-1) = 2 A2 = 1 + 1/2^(1-1) + 1/2^(2-1) = 2.5 A3 = 1 + 1/2^(1-1) + 1/2^(2-1) + 1/2^(3-1)= 2.75

OpenStudy (anonymous):

Adding powers of 5 with decimal points of powers of 10

OpenStudy (amistre64):

An = 1 + 2^0 + 2^-1 + 2^-2 + .... +2^-(n-1)

OpenStudy (amistre64):

that works if we can get the closed form from it which should be 1 + geometric sum

OpenStudy (amistre64):

(1-(1/2)^n)/(1-(1/2)) right?

OpenStudy (amistre64):

that works

OpenStudy (amistre64):

\[A_n=1+\frac{1-\frac{1}{2}^n}{1-\frac{1}{2}}\] \[A_n=1+\frac{1-\frac{1}{2}^n}{\frac{1}{2}}\] \[A_n=1+2(1-\frac{1}{2}^n)\] \[A_n=1+2-2(\frac{1}{2})^n)\] \[A_n=3-2(\frac{1}{2})^n)\]

OpenStudy (amistre64):

hmmm.... its good up to: \[A_n=1+2(1-\frac{1}{2^n})\]

OpenStudy (amistre64):

im assuming we start at A1

OpenStudy (amistre64):

\[A_n=1+2-\frac{2}{2^n}\] \[A_n=3-2^{1-n}\] that one works for me